celticcat:

I got this problem from an old Algebra book and the second part has got me beaten. Part 1 What is the nth term of the sequence 1, 3/2, 5/4, 7/8 ...... Part 2 If the sum of n terms is denoted by s, write down the series obtained by subtracting 1/2 s from s, then find its value in terms of n. Hence find s. I got part 1 to be (2n - 1) / 2^(n-1).

SemiDefinite:

Specialise to find the pattern: \(S_5 = 1 + \frac{3}{2} + \frac{5}{4}+ \frac{7}{8}+ \frac{9}{16}\) \(\frac{S_5}{2} = ~~~~~\frac{1}{2} + \frac{3}{4} + \frac{5}{8} + \frac{7}{ 16} + \frac{9}{32}\) \(~~~\) \(S_5 - \frac{S_5}{2} = 1 + (1 + \frac{1}{2} + \frac{1}{4} + \frac{1}{8} ) - \frac{9}{32}\) Suggests \(\frac{S_n}{2} = 1 + (\text{geometric series}) - \frac{2n-1}{2^{n}}\) Geometric series starts at 1 and has 5 - 1 = 4 terms, constant ratio \(\frac{1}{2}\) Sum of geo series (n-1 terms) is: \( \frac{1 - (\frac{1}{2})^{n-1}}{1 - \frac{1}{2}} = 2 - \frac{1}{2^{n-2}}\) \(\therefore \frac{S_n}{2} = 1 + 2 - \frac{1}{2^{n-2}} - \frac{2n-1}{2^{n}}\) \(= 3 - \frac{3 + 2n}{2^{n}} \) \(\therefore S_n = 6 - \frac{3 + 2n}{2^{n-1}} \) For the inductive proof: \(S_{n+1} = 6 - \frac{3 + 2n}{2^{n-1}} \qquad + \frac{2(n+1)-1}{2^{n - 1 + 1}} \) \( = 6 - \frac{6 + 4n}{2^{n}} + \frac{2n+1}{2^n}\)\) \( = 6 - \frac{5 + 2n}{2^{n}} \) \( = 6 - \frac{3 + 2(n+1)}{2^{n}} \)

celticcat:

Thank you very much. I got the first 3 lines then froze!!

SemiDefinite:

Same!